\begin{thebibliography}{33}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Barbieri et~al.(2015{\natexlab{a}})Barbieri, Ronzano, and
  Saggion}]{Barbieri-IJCAI-2015}
Francesco Barbieri, Francesco Ronzano, and Horacio Saggion. 2015{\natexlab{a}}.
\newblock {Do we criticise (and laugh) in the same way? Automatic detection of
  multi-lingual satirical news in Twitter}.
\newblock In \emph{Proceedings of IJCAI}, pages 1215--1221.

\bibitem[{Barbieri et~al.(2015{\natexlab{b}})Barbieri, Ronzano, and
  Saggion}]{Barbieri-PLN-2015}
Francesco Barbieri, Francesco Ronzano, and Horacio Saggion. 2015{\natexlab{b}}.
\newblock {Is this tweet satirical? A computational approach for satire
  detection in Spanish}.
\newblock \emph{Procesamiento de Lenguaje Natural}, 55:135--142.

\bibitem[{Burfoot and Baldwin(2009)}]{Burfoot-ACL-2009}
Clint Burfoot and Timothy Baldwin. 2009.
\newblock {Automatic Satire Detection: Are You Having a Laugh?}
\newblock In \emph{Proceedings of ACL-IJCNLP}, pages 161--164.

\bibitem[{Butnaru and Ionescu(2019)}]{Butnaru-ACL-2019}
Andrei Butnaru and Radu~Tudor Ionescu. 2019.
\newblock {MOROCO: The Moldavian and Romanian Dialectal Corpus}.
\newblock In \emph{Proceedings of ACL}, pages 688--698.

\bibitem[{De~Sarkar et~al.(2018)De~Sarkar, Yang, and
  Mukherjee}]{DeSarkar-COLING-2018}
Sohan De~Sarkar, Fan Yang, and Arjun Mukherjee. 2018.
\newblock {Attending Sentences to detect Satirical Fake News}.
\newblock In \emph{Proceedings of COLING}, pages 3371--3380.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{Devlin-NAACL-2019}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock {BERT: Pre-training of Deep Bidirectional Transformers for Language
  Understanding}.
\newblock In \emph{Proceedings of NAACL}, pages 4171--4186.

\bibitem[{Dozat(2016)}]{Dozat-ICLR-2016}
Timothy Dozat. 2016.
\newblock {Incorporating Nesterov Momentum into Adam}.
\newblock In \emph{Proceedings of ICLR Workshops}.

\bibitem[{Dumitrescu et~al.(2020)Dumitrescu, Avram, and
  Pyysalo}]{Dumitrescu-EMNLP-2020}
{\c{S}tefan Daniel}~Dumitrescu, Andrei-Marius Avram, and Sampo Pyysalo. 2020.
\newblock {The birth of Romanian BERT}.
\newblock In \emph{Findings of EMNLP}, pages 4324--4328.

\bibitem[{Frain and Wubben(2016)}]{Frain-LREC-2016}
Alice Frain and Sander Wubben. 2016.
\newblock {SatiricLR: a Language Resource of Satirical News Articles}.
\newblock In \emph{Proceedings of LREC}, pages 4137--4140.

\bibitem[{Giudice(2018)}]{Giudice-EVALITA-2018}
Valentino Giudice. 2018.
\newblock {Aspie96 at IronITA (EVALITA 2018): Irony Detection in Italian Tweets
  with Character-Level Convolutional RNN}.
\newblock In \emph{Proceedings of EVALITA}, pages 160--165.

\bibitem[{Goldwasser and Zhang(2016)}]{Goldwasser-TACL-2016}
Dan Goldwasser and Xiao Zhang. 2016.
\newblock {Understanding Satirical Articles Using Common-Sense}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  4:537--549.

\bibitem[{Ionescu and Chifu(2021)}]{Ionescu-Arxiv-2021}
Radu~Tudor Ionescu and Adrian~Gabriel Chifu. 2021.
\newblock Fresada: A french satire data set for cross-domain satire detection.
\newblock \emph{arXiv preprint arXiv:2104.04828}.

\bibitem[{Jia et~al.(2019)Jia, Deng, Min, and Liu}]{Jia2019}
Xiuyi Jia, Zhao Deng, Fan Min, and Dun Liu. 2019.
\newblock {Three-way decisions based feature fusion for Chinese irony
  detection}.
\newblock \emph{International Journal of Approximate Reasoning}, 113:324--335.

\bibitem[{Karoui et~al.(2017)Karoui, Zitoune, and Moriceau}]{Karoui2017}
Jihen Karoui, Farah Zitoune, and Véronique Moriceau. 2017.
\newblock {SOUKHRIA: Towards an Irony Detection System for Arabic in Social
  Media}.
\newblock In \emph{Proceedings of ACLing}, volume 117, pages 161--168.

\bibitem[{Klambauer et~al.(2017)Klambauer, Unterthiner, Mayr, and
  Hochreiter}]{Klambauer-NIPS-2017}
G\"{u}nter Klambauer, Thomas Unterthiner, Andreas Mayr, and Sepp Hochreiter.
  2017.
\newblock {Self-Normalizing Neural Networks}.
\newblock In \emph{Proceedings of NIPS}, pages 972--981.

\bibitem[{Li et~al.(2020)Li, Levi, Hosseini, and Broniatowski}]{Li-NLP4IF-2020}
Lily Li, Or~Levi, Pedram Hosseini, and David Broniatowski. 2020.
\newblock {A Multi-Modal Method for Satire Detection using Textual and Visual
  Cues}.
\newblock In \emph{Proceedings of NLP4IF}, pages 33--38.

\bibitem[{Liebrecht et~al.(2013)Liebrecht, Kunneman, and van~den
  Bosch}]{Liebrecht2013}
Christine Liebrecht, Florian Kunneman, and Antal van~den Bosch. 2013.
\newblock The perfect solution for detecting sarcasm in tweets {\#}not.
\newblock In \emph{Proceedings of WASSA}, pages 29--37.

\bibitem[{Liu et~al.(2019)Liu, Shabani, Balet, and Sokhn}]{Liu-ICCCN-2019}
Zhan Liu, Shaban Shabani, Nicole~Glassey Balet, and Maria Sokhn. 2019.
\newblock {Detection of Satiric News on Social Media: Analysis of the
  Phenomenon with a French Dataset}.
\newblock In \emph{Proceedings of ICCCN}, pages 1--6.

\bibitem[{Loshchilov and Hutter(2019)}]{Loshchilov-ICLR-2019}
Ilya Loshchilov and Frank Hutter. 2019.
\newblock {Decoupled Weight Decay Regularization}.
\newblock In \emph{Proceedings of ICLR}.

\bibitem[{McClennen and Maisel(2014)}]{McClennen-Springer-2014}
Sophia~A. McClennen and Remy~M. Maisel. 2014.
\newblock {I'm Not Laughing at You, I'm Laughing With You: How to Stop Worrying
  and Love the Laughter}.
\newblock In \emph{Is Satire Saving Our Nation? Mockery and American Politics},
  pages 189--201. Springer.

\bibitem[{McHardy et~al.(2019)McHardy, Adel, and Klinger}]{McHardy-NAACL-2019}
Robert McHardy, Heike Adel, and Roman Klinger. 2019.
\newblock {Adversarial Training for Satire Detection: Controlling for
  Confounding Variables}.
\newblock In \emph{Proceedings of NAACL}, pages 660--665.

\bibitem[{Meel and Vishwakarma(2019)}]{Meel-ESA-2019}
Priyanka Meel and Dinesh~Kumar Vishwakarma. 2019.
\newblock {Fake news, rumor, information pollution in social media and web: A
  contemporary survey of state-of-the-arts, challenges and opportunities}.
\newblock \emph{Expert Systems with Applications}, 153:112986.

\bibitem[{P{\'e}rez-Rosas et~al.(2018)P{\'e}rez-Rosas, Kleinberg, Lefevre, and
  Mihalcea}]{Perez-COLING-2018}
Ver{\'o}nica P{\'e}rez-Rosas, Bennett Kleinberg, Alexandra Lefevre, and Rada
  Mihalcea. 2018.
\newblock {Automatic Detection of Fake News}.
\newblock In \emph{Proceedings of COLING}, pages 3391--3401.

\bibitem[{Peters and Broersma(2013)}]{Peters-Routledge-2013}
Chris Peters and Marcel~Jeroen Broersma. 2013.
\newblock \emph{{Rethinking Journalism: Trust and Participation in a
  Transformed News Landscape}}.
\newblock Routledge.

\bibitem[{Ravi and Ravi(2017)}]{Ravi-KBS-2017}
Kumar Ravi and Vadlamani Ravi. 2017.
\newblock A novel automatic satire and irony detection using ensembled feature
  selection and data mining.
\newblock \emph{Knowledge-Based Systems}, 120:15--33.

\bibitem[{Rubin et~al.(2016)Rubin, Conroy, Chen, and
  Cornwell}]{Rubin-CADD-2016}
Victoria Rubin, Niall Conroy, Yimin Chen, and Sarah Cornwell. 2016.
\newblock {Fake News or Truth? Using Satirical Cues to Detect Potentially
  Misleading News}.
\newblock In \emph{Proceedings of CADD}, pages 7--17.

\bibitem[{Saadany et~al.(2020)Saadany, Orasan, and Mohamed}]{Saadany-RDSM-2020}
Hadeel Saadany, Constantin Orasan, and Emad Mohamed. 2020.
\newblock {Fake or Real? A Study of Arabic Satirical Fake News}.
\newblock In \emph{Proceedings of RDSM}, pages 70--80.

\bibitem[{Selvaraju et~al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and
  Batra}]{Selvaraju-ICCV-2017}
Ramprasaath~R. Selvaraju, Michael Cogswell, Abhishek Das, Ramakrishna Vedantam,
  Devi Parikh, and Dhruv Batra. 2017.
\newblock {Grad-CAM: Visual Explanations from Deep Networks via Gradient-based
  Localization}.
\newblock In \emph{Proceedings of ICCV}, pages 618--626.

\bibitem[{Sharma et~al.(2019)Sharma, Qian, Jiang, Ruchansky, Zhang, and
  Liu}]{Sharma-TIST-2019}
Karishma Sharma, Feng Qian, He~Jiang, Natali Ruchansky, Ming Zhang, and Yan
  Liu. 2019.
\newblock {Combating fake news: A survey on identification and mitigation
  techniques}.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology},
  10(3):1--42.

\bibitem[{To{\c{c}}o{\u{g}}lu and Onan(2019)}]{Tocoglu-ID-2019}
Mansur~Alp To{\c{c}}o{\u{g}}lu and Aytu{\u{g}} Onan. 2019.
\newblock {Satire Detection in Turkish News Articles: A Machine Learning
  Approach}.
\newblock In \emph{Proceedings of Innovate-Data}, pages 107--117.

\bibitem[{Yang et~al.(2017)Yang, Mukherjee, and Dragut}]{Yang-EMNLP-2017}
Fan Yang, Arjun Mukherjee, and Eduard Dragut. 2017.
\newblock {Satirical News Detection and Analysis using Attention Mechanism and
  Linguistic Features}.
\newblock In \emph{Proceedings EMNLP}, pages 1979--1989.

\bibitem[{Zhang et~al.(2015)Zhang, Zhao, and LeCun}]{Zhang-NIPS-2015}
Xiang Zhang, Junbo Zhao, and Yann LeCun. 2015.
\newblock {Character-level Convolutional Networks for Text Classification}.
\newblock In \emph{Proceedings of NIPS}, pages 649--657.

\bibitem[{Zhang et~al.(2020)Zhang, Yang, Dragut, and
  Mukherjee}]{Zhang-ArXiv-2020}
Yigeng Zhang, Fan Yang, Eduard~Constantin Dragut, and Arjun Mukherjee. 2020.
\newblock {Birds of a Feather Flock Together: Satirical News Detection via
  Language Model Differentiation}.
\newblock \emph{arXiv preprint arXiv:2007.02164}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2021}
\usepackage{times}
\usepackage{latexsym}
\usepackage{float}
\restylefloat{table}

\usepackage{microtype}
\usepackage{graphicx}
\usepackage{xcolor}
\usepackage{epsfig}
\usepackage{url}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{multirow}

\aclfinalcopy 
\def\aclpaperid{1170} 

\title{\textcolor{blue}{Sa}\textcolor{yellow}{Ro}\textcolor{red}{Co}: Detecting \textcolor{blue}{Sa}tire in a Novel \textcolor{yellow}{Ro}manian \textcolor{red}{Co}rpus of News Articles}

\author{Ana-Cristina Rogoz, Mihaela G\u{a}man, Radu Tudor Ionescu* \\
  University of Bucharest\\
  14 Academiei Street, Bucharest, Romania \\
  \texttt{*raducu.ionescu@gmail.com}}

\date{}

\begin{document}
\maketitle

\begin{abstract}
In this work, we introduce a corpus for satire detection in Romanian news. We gathered 55,608 public news articles from multiple real and satirical news sources, composing one of the largest corpora for satire detection regardless of language and the only one for the Romanian language. We provide an official split of the text samples, such that training news articles belong to different sources than test news articles, thus ensuring that models do not achieve high performance simply due to overfitting. We conduct experiments with two state-of-the-art deep neural models, 
resulting in a set of strong baselines for our novel corpus. Our results show that the machine-level accuracy for satire detection in Romanian is quite low (under $73\%$ on the test set) compared to the human-level accuracy ($87\%$), leaving enough room for improvement in future research.
\end{abstract}

\section{Introduction}
\label{sec_intro}

According to its definition in the Cambridge Dictionary, satire is ``a humorous way of criticizing people or ideas''
. News satire employs this mechanism in the form of seemingly legitimate journalistic reporting, with the intention of ridiculing public figures, politics or contemporary events \cite{McClennen-Springer-2014, Peters-Routledge-2013,Rubin-CADD-2016}. Although the articles pertaining to this genre contain fictionalized stories, the intent is not to mislead the public into thinking that the discussed subjects are real. On the contrary, satirical news articles are supposed to reveal their nature by the writing style and comedic devices employed, such as irony, parody or exaggeration. Thus, the intention behind the writing differentiates satirical news \cite{Rubin-CADD-2016} from fake news \cite{Meel-ESA-2019,Perez-COLING-2018,Sharma-TIST-2019}. However, in some rare cases, the real intent might be deeply buried in the complex irony and subtleties of news satire \cite{Barbieri-IJCAI-2015}, which has the effect of fiction being deemed as factual \cite{Zhang-ArXiv-2020}. Even so, there is a clear distinction between satirical and fake news. In fake news, the intent is to deceive the readers in thinking that the news is real, while presenting fake facts to influence the readers' opinion. Since our study is focused on satire detection, we consider discussing research on fake news detection as being out of our scope. 
At the same time, we acknowledge the growing importance of detecting fake news and the fact that an accurate differentiation of satirical from legitimate journalistic reports might be seen as a starting point in controlling the spread of deceptive news \cite{DeSarkar-COLING-2018}.

\begin{table*}[t]
\begin{center}
\footnotesize
\begin{tabular}{|l|l|r|r|r|}
\hline
\multirow{2}{*}{Data Set}					& \multirow{2}{*}{Language}      & \multicolumn{3}{|c|}{\#articles}  \\
\cline{3-5}
     						&               & Regular       & Satirical	        &  Total\\
\hline
\hline
\cite{Burfoot-ACL-2009}     & English       & 4,000			& 233		        & 4,233 \\
\cite{Frain-LREC-2016}      & English       & 1,705			& 1,706		        & 3,411 \\
\cite{Goldwasser-TACL-2016} & English       & 10,921		& 1,225		        & 12,146 \\
\cite{Ionescu-Arxiv-2021}   & French        & 5,648         & 5,922             & 11,570 \\
\cite{Li-NLP4IF-2020}       & English       & 6,000			& 4,000		        & 10,000 \\
\cite{Liu-ICCCN-2019}       & French        & 2,841         & 2,841             & 5,682 \\
\cite{McHardy-NAACL-2019}   & German        & 320,219 		& 9,643 		    & 329,862 \\
\cite{Ravi-KBS-2017}		& English       & 1,272			& 393		        & 1,665 \\
\cite{Saadany-RDSM-2020}    & Arabic        & 3,185			& 3,710		        & 6,895 \\
\cite{Tocoglu-ID-2019}      & Turkish       & 1,000			& 1,000		        & 2,000 \\
\cite{Yang-EMNLP-2017}      & English       & 168,780		& 16,249		    & 185,029 \\
\hline
SaRoCo (ours)				& Romanian      & 27,980		& 27,628		    & 55,608  \\
\hline
\end{tabular}
\end{center}
\caption{Number of regular and satirical news articles in existing corpora versus SaRoCo.}
\label{tab_corpora}
\end{table*}


\begin{table*}[t]
\begin{center}
\footnotesize
\begin{tabular}{|l|r|r|r|r|r|r|}
\hline
\multirow{2}{*}{Set} 						& \multicolumn{2}{|c|}{Regular}     & \multicolumn{2}{|c|}{Satirical}   & \multicolumn{2}{|c|}{Total}\\
\cline{2-7}
     						& \#articles		& \#tokens	        & \#articles		& \#tokens          & \#articles		& \#tokens\\
\hline
\hline
Training				    & 18,000 		& 8,174,820         & 17,949 		& 11,147,169         & 35,949        & 19,321,989\\
Validation					& 4,986			& 2,707,026           & 4,878		    & 3,030,055           & 9,864         & 5,737,081\\
Test					    & 4,994			& 2,124,346           & 4,801		    & 1,468,199           & 9,795         & 3,592,545\\
\hline
Total						& 27,980		& 13,006,192         & 27,628		& 15,645,423         & 55,608        & 28,651,615\\
\hline
\end{tabular}
\end{center}
\caption{Number of samples (\#articles) and number of tokens (\#tokens) for each subset in SaRoCo.}
\label{tab_SaRoCo}
\end{table*}

Satire detection is an important task that could be addressed prior to the development of conversational systems and robots that interact with humans. Certainly, the importance of understanding satirical (funny, ridiculous or ironical) text becomes obvious when we consider a scenario in which a robot performs a dangerous action because it takes a satirical comment of the user too literally. Given the relevance of the task for the natural language processing community, satire detection has already been investigated in several well-studied languages such as Arabic \cite{Saadany-RDSM-2020}, English \cite{Burfoot-ACL-2009,DeSarkar-COLING-2018,Goldwasser-TACL-2016,Yang-EMNLP-2017}, French \cite{Ionescu-Arxiv-2021,Liu-ICCCN-2019}, German \cite{McHardy-NAACL-2019}, Spanish \cite{Barbieri-PLN-2015} and Turkish \cite{Tocoglu-ID-2019}. Through the definition of satire, the satire detection task is tightly connected to irony and sarcasm detection. These tasks strengthen or broaden the language variety with languages such as Arabic \cite{Karoui2017}, Chinese \cite{Jia2019}, Dutch \cite{Liebrecht2013} and Italian \cite{Giudice-EVALITA-2018}. 

In this work, we introduce SaRoCo\footnote{\url{https://github.com/MihaelaGaman/SaRoCo}}, the \textbf{Sa}tire detection \textbf{Ro}manian \textbf{Co}rpus, which comprises 55,608 news articles collected from various sources. To the best of our knowledge, this is the first and only data set for the study of Romanian satirical news. Furthermore, SaRoCo is also one of the largest data sets for satirical news detection, being surpassed only two corpora, one for English \cite{Yang-EMNLP-2017} and one for German \cite{McHardy-NAACL-2019}. However, our corpus contains the largest collection of satirical news articles (over 27,000). These facts are confirmed by the comparative statistics presented in Table~\ref{tab_corpora}.

Along with the novel data set, we include two strong deep learning methods to be used as baselines in future works. The first method is based on low-level features learned by a character-level convolutional neural network \cite{Zhang-NIPS-2015}, while the second method employs high-level semantic features learned by the Romanian version of BERT \cite{Dumitrescu-EMNLP-2020}. The gap between the human-level performance and that of the deep learning baselines 
indicates that there is enough room for improvement left for future studies. We make our corpus and baselines available online for nonprofit educational and research purposes, under an open-source noncommercial license agreement.



\section{Corpus}
\label{sec:benchmark}

SaRoCo gathers both satirical and non-satirical news from some of the most popular Romanian news websites. The collected news samples were found in the public web domain, i.e.~access is provided for free without requiring any subscription to the publication sources. The entire corpus consists of 55,608 samples (27,628 satirical samples and 27,980 non-satirical samples), having more than 28 million tokens in total, as illustrated in Table \ref{tab_SaRoCo}. Each sample is composed of a title (headline), a body and a corresponding label (satirical or non-satirical). As shown in Table~\ref{tab_avg_stats}, an article has around $515.24$ tokens on average, with an average of $24.97$ tokens for the headline. We underline that the labels are automatically determined, based on the fact that a publication source publishes either regular or satirical news, but not both.

\begin{table}[!t]
\footnotesize
\begin{center}
 \begin{tabular}{|l | r |} 
 \hline
 Sample Part    & Average \#tokens\\
 \hline 
 \hline
 Title          & 24.97 \\ 
 Full Articles  & 515.24 \\
 \hline
\end{tabular}
\end{center}
\caption{Average number of tokens in full news articles and titles from SaRoCo.}
\label{tab_avg_stats}
\end{table}

We provide an official split for our corpus, such that all future studies will use the same training, validation and test sets, easing the direct comparison with prior results. Following \newcite{McHardy-NAACL-2019}, we use disjoint sources for training, validation and test, ensuring that models do not achieve high performance by learning author styles or topic biases particular to certain news websites. While crawling the public news articles, we selected the same topics (culture, economy, politics, social, sports, tech) and the same time frame (between 2011 and 2020) for all news sources to control for potential biases induced by uneven topic or time distributions across the satirical and non-satirical genres.

\begin{table}[t]
\setlength\tabcolsep{4.8pt}
\centering
\footnotesize
\begin{tabular}{|p{0.16\linewidth}|p{0.34\linewidth}|p{0.35\linewidth}|}
\hline
Category &  Example & Translation \\
\hline 
\hline
\multirow{13}{*}{Regular}            & \textit{``Tragedie \^{i}n zi de s\u{a}rb\u{a}toare''} & \textit{``Tragedy during celebration day''} \\
\cline{2-3}
            & \textit{``Demisia lui \$NE\$ \$NE\$ se am\^{a}n\u{a}''} & \textit{``\$NE\$ \$NE\$'s resignation is post-poned''} \\
\cline{2-3}
            & \textit{``Premierul bulgar \$NE\$ \$NE\$ are \$NE\$''} & \textit{``Bulgarian prime-minister \$NE\$ \$NE\$ has \$NE\$''} \\
\cline{2-3}
            & \textit{``A murit actorul \$NE\$ \$NE\$''} & \textit{``The actor \$NE\$ \$NE\$ died''} \\
\cline{2-3}
            & \textit{``Metroul din \$NE\$ \$NE\$ se deschide azi''} & \textit{``Subway to \$NE\$ \$NE\$ opens up today''} \\
\hline
\multirow{15}{*}{Satirical}        & \textit{``Comedia cu p\u{a}l\u{a}rioar\u{a} de staniol''} &  \textit{``Comedy with little tinfoil hat''} \\
\cline{2-3}
        & \textit{``10 restric\c{t}ii dure pe care \$NE\$ le preg\u{a}te\c{s}te pe ascuns''} & \textit{``10 harsh restrictions that \$NE\$ is planning in secrecy''} \\
\cline{2-3}
 & \textit{``C\^{a}\c{t}i pokemoni ai prins azi?''} & \textit{``How many pokemons did you catch today?''}\\
\cline{2-3}        
        & \textit{``Biserica \$NE\$ lanseaz\u{a} apa sfin\c{t}it\u{a} cu arom\u{a}''} & \textit{``The \$NE\$ Church launches flavored holy water''}\\
\cline{2-3}
        & \textit{``Dragostea \^{i}n vremea sclerozei''} & \textit{``Love in the time of sclerosis''} \\
\hline
\end{tabular}
\caption{Examples of news headlines from SaRoCo.}
\label{tab_examples_news_headlines}
\end{table}

\begin{table*}[!t]
\setlength\tabcolsep{4.2pt}
\begin{center}
\footnotesize
\begin{tabular}{|l|c|c|c|c|c|c|c|c|c|c|c|c|c|c|c|c|}
\hline 
\multirow{3}{*}{Method} & \multicolumn{6}{|c|}{Validation} & \multicolumn{6}{|c|}{Test} \\
\cline{2-13}
 						&  \multirow{2}{*}{Acc.} & Macro & \multicolumn{2}{|c|}{Satirical} & \multicolumn{2}{|c|}{Regular} & \multirow{2}{*}{Acc.} & Macro & \multicolumn{2}{|c|}{Satirical} & \multicolumn{2}{|c|}{Regular} \\
\cline{4-7} \cline{10-13}
&  & $F_1$ & Prec. & Rec. & Prec. & Rec. & & $F_1$ & Prec. & Rec. & Prec. & Rec.\\
\hline
\hline
Ro-BERT  & 0.8241	& 0.8160	& 0.9260	& 0.6991	& 0.7633	& 0.9462 &	0.7300	& 0.7150 &	0.8750 &	0.5250 &	0.6700 &	0.9250 \\
Char-CNN & 0.7342 & 0.7475 & 0.8023 & 0.6138 & 0.6928 & 0.8520 & 0.6966 & 0.7109 & 0.7612 & 0.5551 & 0.6606 & 0.8326 \\
\hline
\end{tabular}
\end{center}
\caption{Validation and test results of the character-level CNN and the fine-tuned Ro-BERT applied on SaRoCo.}
\label{tab_results_full}
\end{table*}

After crawling satirical and non-satirical news samples, our first aim was to prevent discrimination based on named entities. The satirical character of an article should be inferred from the language use rather than specific clues, such as named entities. For example, certain sources of news satire show preference towards mocking politicians from a specific political party, and an automated system might erroneously label a news article about a member of the respective party as satirical simply based on the presence of the named entity. Furthermore, we even noticed that some Romanian politicians have certain mocking nicknames assigned in satirical news. In order to eliminate named entities, we followed a similar approach as the one used for the MOROCO \cite{Butnaru-ACL-2019} data set. Thus, all the identified named entities are replaced with the special token \$NE\$.
Besides eliminating named entities, we also substituted all whitespace characters with space and replaced multiple consecutive spaces with a single space. A set of processed satirical and regular headlines are shown in Table~\ref{tab_examples_news_headlines}.






\section{Baselines}

\noindent
\textbf{Fine-tuned Ro-BERT.}
Our first baseline consists of a fine-tuned Romanian BERT \cite{Dumitrescu-EMNLP-2020}, which follows the same transformer-based model architecture as the original BERT \cite{Devlin-NAACL-2019}. According to \newcite{Dumitrescu-EMNLP-2020}, the Romanian BERT (Ro-BERT) attains better results than the multilingual BERT on a range of tasks. We therefore assume that the Romanian BERT should represent a stronger baseline for our Romanian corpus.

We use the Ro-BERT encoder to encode each text sequence into a list of token IDs. The tokens are further processed by the model, obtaining the corresponding 768-dimensional embeddings. At this point, we add a global average pooling layer to obtain a Continuous Bag-of-Words (CBOW) representation for each sequence of text, followed by a Softmax output layer with two neural units, each predicting the probability for one category, either non-satirical or satirical. To obtain the final class label for a text sample, we apply \emph{argmax} on the two probabilities. We fine-tune the whole model for 10 epochs on mini-batches of 32 samples, using the Adam with decoupled weight decay (AdamW) optimizer \cite{Loshchilov-ICLR-2019}, with a learning rate of $10^{-7}$ and the default value for $\epsilon$.

\noindent
\textbf{Character-level CNN.}
The second baseline model considered in the experiments is a Convolutional Neural Network (CNN) that operates at the character level \cite{Zhang-NIPS-2015}. We set the input size to 1,000 characters. 
After the input layer, we add an embedding layer to encode each character into a vector of $128$ components. The optimal architecture for the task at hand proved to be composed of three convolutional (conv) blocks, each having a conv layer with $64$ filters applied at stride $1$, followed by Scaled Exponential Linear Unit (SELU) activation. From the first block to the third block, the convolutional kernel sizes are $5$, $3$ and $1$, respectively. Max-pooling with a filter size of $3$ is applied after each conv layer. After each conv block, we insert a Squeeze-and-Excitation block with the reduction ratio set to $r = 64$, following \newcite{Butnaru-ACL-2019}. To prevent overfitting, we use batch normalization and Alpha Dropout \cite{Klambauer-NIPS-2017} with a dropout rate of 
$0.5$. The final prediction layer is composed of two neural units, one for each class (i.e.~legitimate and satirical), with Softmax activation. We use the Nesterov-accelerated Adaptive Moment Estimation (Nadam) optimizer \cite{Dozat-ICLR-2016} with a learning rate of $2\cdot10^{-4}$, training the network for $50$ epochs on mini-batches of $128$ samples.

\section{Experiments}

\noindent
\textbf{Evaluation.}
We conducted binary classification experiments on SaRoCo, predicting if a given piece of text is either satirical or non-satirical. As evaluation metrics, we employ the precision and recall for each of the two classes. We also combine these scores through the macro $F_1$ and micro $F_1$ (accuracy) measures.

\noindent
\textbf{Results.}
In Table~\ref{tab_results_full}, we present the results of the two baselines on the SaRoCo validation and test sets. We observe that both models tend to have higher precision scores in detecting satire than in detecting regular news. The trade-off between precision and recall is skewed towards higher recall for the non-satirical news class. Since both models share the same behavior, we conjecture that the behavior is rather caused by the particularities of the satire detection task.

\begin{table}[!t]
\setlength\tabcolsep{4.8pt}
\centering
\footnotesize
\begin{tabular}{|p{0.2\linewidth}|p{0.31\linewidth}|p{0.33\linewidth}|}
\hline
{Category} &  Example & Translation \\
\hline 
\hline
Slang & \textit{``cel mai marf\u{a} serial din lume''} & \textit{``the dopest TV show in the world''} \\
                     & \textit{``cocalar''} & \textit{``douche''} \\
\hline
Insult & \textit{``odiosul primar''} & \textit{``the odious mayor''}\\
                      & \textit{``bunicu\c{t} retardat''} & \textit{``retarded grandpa''}\\
                      & \textit{``dugongul \u{a}la slinos de la sectorul 4''} & \textit{``that slender dugong in the 4th sector''}\\
\hline
Repetition & \textit{``Mii de gunoaie care las\u{a} gunoaie au remarcat c\u{a} [...] plajele [...] s-au umplut de gunoaie, l\u{a}sate [...] de gunoaiele care au venit \^{i}naintea lor''} & \textit{``Thousands of scums who leave garbage noticed that [...] beaches [...] got full of garbage, left behind [...] by the scums who were there before them''}\\
\hline
Exaggeration    & \textit{``Ne-am s\u{a}turat!''}                   & \textit{``We're sick of it!''}\\
Exclamation    & \textit{``Ru\c{s}ine s\u{a} le fie!''}                & \textit{``Shame on them!''}\\
Irony           & \textit{``Chiar nu suntem o na\c{t}ie de ho\c{t}i!''} & \textit{``We're totally not a nation of thieves!''}\\
\hline
Popular & \textit{``a s\u{a}rit calul''}     & \textit{``went overboard''}\\
Saying & \textit{``a f\u{a}cut-o de oaie''} & \textit{``messed up''}\\
                      & \textit{``minte de g\u{a}in\u{a}''}    & \textit{``bird brain''}\\
\hline
\end{tabular}
\caption{Examples of predictive patterns of satire learned by the character-level CNN.}
\label{tab_most_predictive_patterns_satire}
\end{table}


\begin{table}[!t]
\setlength\tabcolsep{4.8pt}
\centering
\footnotesize
\begin{tabular}{|p{0.2\linewidth}|p{0.31\linewidth}|p{0.33\linewidth}|}
\hline
Category &  Example & Translation \\
\hline 
\hline
Stats & \textit{``Importurile au sc\u{a}zut cu $2.1\%$ [...] pentru o cre\c{s}tere de $0.1\%$ \c{s}i prelungirea sc\u{a}derii de $1.4\%$ din iulie.''} & \textit{``Imports decreased by $2.1\%$ [...] for an increase of $0.1\%$ and the prolongation of the decrease of $1.4\%$ since July.''} \\
\hline
Legal terms   & \textit{``asasinat''} & \textit{``assasinated''} \\
                        & \textit{``l-au denun\c{t}at pe autorul atacului''} & \textit{``denounced the perpetrator''} \\
\hline
Weather    & \textit{``temperatura \^{i}n timpul nop\c{t}ii a sc\u{a}zut''} & \textit{``the temperature has dropped during the night''} \\
\hline
Political terms     & \textit{``scrutinul preziden\c{t}ial''} & \textit{``presidential election''} \\
                    & \textit{``prefectura informeaz\u{a} c\u{a}''} & \textit{``prefecture informs that''} \\
\hline
\end{tabular}
\caption{Examples of predictive patterns of legitimate news learned by the character-level CNN.}
\label{tab_most_predictive_patterns_legitimate}

\end{table}

\noindent
\textbf{Discriminative feature analysis.}
We analyze the discriminative features learned by the character-level CNN, which is one of the proposed baseline systems for satire detection. We opted for the character-level CNN in favor of the fine-tuned BERT, as the former method allows us to visualize discriminative features using Grad-CAM~\cite{Selvaraju-ICCV-2017}, a technique that was initially used to explain decisions of CNNs applied on images. We adapted this technique for the character-level CNN, then extracted and analyzed the most predictive patterns in SaRoCo. The motivation behind this was to validate that the network's decisions are not based on some biases that escaped our data collection and cleaning process. 

\begin{table*}[!t]
\centering
\footnotesize
\begin{tabular}{|l|c|c|c|c|c|c|}
\hline
\multirow{2}{*}{Method} & \multirow{2}{*}{Acc.} & Macro & \multicolumn{2}{|c|}{Satirical} & \multicolumn{2}{|c|}{Regular} \\
\cline{4-7}
& & $F_1$ & Prec. & Rec. & Prec. & Rec.\\
\hline
\hline 
Ro-BERT & 0.6800	& 0.6750 &	0.7800 &	0.5100 &	0.6350 &	0.8550 \\
Char-CNN & 0.6500 &  0.6510 & 0.6389 & 0.6900 & 0.6630 & 0.6100 \\
\hline
Humans & 0.8735 & 0.8711 & 0.9416 & 0.7970 & 0.8332 & 0.9500 \\
\hline
\end{tabular}
\caption{Averaged performance of ten human annotators versus deep learning baselines on 200 news headlines from SaRoCo. 
}
\label{tab_human_eval}
\end{table*}

In Tables~\ref{tab_most_predictive_patterns_satire} and~\ref{tab_most_predictive_patterns_legitimate}, we present a few examples of interesting patterns considered relevant for predicting satire versus regular news, respectively. A broad range of constructions covering a great variety of styles and significant words are underlined via Grad-CAM in the satirical news samples. The network seems to pick up obvious clues such as slang, insults and popular sayings rather than more subtle indicatives of satire, including irony or exaggeration. 
At the same time, for the real news in SaRoCo, there are fewer categories of predictive patterns. In general, the CNN deems formal, standard news expressions as relevant for regular news. These patterns vary across topics and domains. The CNN also finds that the presence of numbers and statistical clues is indicative for non-satirical content, which is consistent with the observations of \newcite{Yang-EMNLP-2017}. Our analysis reveals that the discriminative features are appropriate for satire detection, showing that our corpus is indeed suitable for the considered task.

\noindent
\textbf{Deep models versus humans.}
Given 100 satirical and 100 non-satirical news headlines (titles) randomly sampled from the SaRoCo test set, we asked ten Romanian human annotators to label each sample as satirical or non-satirical. We evaluated the deep learning methods on the same subset of 200 samples, reporting the results in Table~\ref{tab_human_eval}. First, we observe that humans have a similar bias as the deep learning models. Indeed, for both humans and models, the trade-off between precision and recall is skewed towards higher precision for the satirical class and higher recall for the non-satirical class. We believe this is linked to the way people and machines make a decision. Humans look for patterns of satire in order to label a sample as satire. If a satire-specific pattern is not identified, the respective sample is labeled as regular, increasing the recall for the non-satirical class. Although humans and machine seem to share the same way of thinking, there is a considerable performance gap in satire detection between humans and machines. Indeed, the average accuracy of our ten human annotators is around $87\%$, while the state-of-the-art deep learning models do not surpass $68\%$ on the same news headlines. Even on full news articles (see Table~\ref{tab_results_full}), the models barely reach an accuracy of $73\%$ on the test set. Hence, we conclude there is a significant performance gap between humans and machines, leaving enough room for exploration in future work on Romanian satire detection.

We would like to emphasize that our human evaluation was performed by casual news readers, and the samples were shown after named entity removal, thus having a fair comparison with the AI models. We underline that named entity removal makes the task more challenging, even for humans.

\section{Conclusion}

In this work, we presented SaRoCo, a novel data set containing satirical and non-satirical news samples. To the best of our knowledge, SaRoCo is the only corpus for Romanian satire detection and one of the largest corpora regardless of language. We trained two state-of-the-art neural models as baselines for future research on our novel corpus. We also compared the performance of the neural models with the averaged performance of ten human annotators, showing that the neural models lag far behind the human-level performance. Our discriminative feature analysis confirms the limitations of state-of-the-art neural models in detecting satire. Although we selected a set of strong models from the recent literature as baselines for SaRoCo, significant future research is necessary to close the gap with respect to the human-level satire detection performance. Designing models to pick up irony or exaggerations could pave the way towards closing this gap in future work.

\section*{Acknowledgments}

The authors thank reviewers for their useful remarks. 
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI – UEFISCDI, project number PN-III-P1-1.1-TE-2019-0235, within PNCDI III. This article has also benefited from the support of the Romanian Young Academy, which is funded by Stiftung Mercator and the Alexander von Humboldt Foundation for the period 2020-2022.



\bibliographystyle{acl_natbib}
\bibliography{acl2021-arxiv}


\end{document}